If the length of a conductor becomes double calculate the value of new resistance.

Grade:12

2 Answers

jogesh
5 years ago
And:-- Let original length of conductor = l. Specific resistance =p Area of cross section of conductor = A. Then resistance R= p×l/A ••••••••••• (1). Now. When length becomes new length of conductor (l*) = 2l specific resistance ( remains same because it is same for a kind of conductor not change with shape and size of conductor ) = p now we have to find new area of conductor we know that volume will not change .•. Volume original conductor = volume change conductor = V= V = length × cross sectional Area l×A= 2l× A* ( A*= Area of changed conductor ) Then. A*= A/2. ••••••••• (2). Now; new resistance R* = p×l*/ A* = 2p× 2l / A= 4pl/A ( because l* = 2l and A* A/2). And we know that pl/A= R then R* = 4(pl/A)= 4R
Khimraj
5 years ago
Resistance of a conductor is direct